OpenStudy (anonymous):

In this problem, you will apply kinematic equations to a jumping flea. Take the magnitude of free-fall acceleration to be 9.80m/s2 . Ignore air resistance. A flea jumps straight up to a maximum height of 0.460m . What is its initial velocity v0 as it leaves the ground?

OpenStudy (souvik):

\[v^2=u^2-2gh\] where v=velocity at highest point=0 u=initial velocity g=9.8 m/s^2 and h=height